# High-resolution building and road detection from Sentinel-2

### v2

@@ Abstract
Mapping buildings and roads automatically with remote sensing typically requires high-resolution imagery, which is expensive to obtain and often sparsely available. In this work we demonstrate how multiple 10 m resolution Sentinel-2 images can be used to generate 50 cm resolution building and road segmentation masks. This is done by training a ‘student’ model with access to Sentinel-2 images to reproduce the predictions of a ‘teacher’ model which has access to corresponding high-resolution imagery. While the predictions do not have all the fine detail of the teacher model, we find that we are able to retain much of the performance: for building segmentation we achieve 78.3% mIoU, compared to the high-resolution teacher model accuracy of 85.3% mIoU. We also describe a related method for counting individual buildings in a Sentinel-2 patch which achieves $R^{2}=0.91$ against true counts. This work opens up new possibilities for using freely available Sentinel-2 imagery for a range of tasks that previously could only be done with high-resolution satellite imagery.
@@ 1 Introduction
Figure 3: Estimation of the number of buildings in a tile, based on predicting building centroids (left: high resolution image for comparison, centre: top of Sentinel-2 stack input to the model; right: predicted centroid mask). This method can obtain $R^{2}=0.91$ with respect to true counts even though individual buildings cannot be discerned in the source imagery.
In our work, we attempt to extend the limit of fine detail that can be recreated from a set of Sentinel-2 images, assessing the quality of building and road presence predictions made at 50 cm resolution. To do this, we use a teacher model with access to 50 cm resolution imagery to create training labels for a large dataset of worldwide imagery. We then train a student model to reconstruct these labels given only a stack of Sentinel-2 images from the corresponding places and times (see Figure 1). We find that we are able to retain much – though not all – of the accuracy of the high-resolution teacher model: our Sentinel-2-based building segmentation has 78.3% mIoU, compared to the high-resolution-based teacher performance of 85.3% mIoU. We found that this accuracy level was equivalent to what could be achieved by a high-resolution model using a single frame of 4 m resolution imagery (see Table 5), though also noting through visual inspection that the segmentation quality sometimes far exceeds that (see Figure 2 and Section 10).
Finally, we describe a method using our framework for counting the number of individual buildings in a patch, based on predicting the locations of building centroids. We can obtain $R^{2}=0.91$ compared to true building counts, again capturing much of the performance of the high-resolution teacher model ($R^{2}=0.95$). The predicted counts are surprisingly accurate even when buildings are very small relative to the raw Sentinel-2 resolution, or close together. Some examples are shown in Figures 3 and 15.
This work extends the range of analysis tasks that can be carried out with freely available Sentinel-2 data. As well as using it to improve the Open Buildings dataset [^1] https://sites.research.google/open-buildings, we plan to make model resources available for social benefit purposes in the near future.
@@ 3 Training setup
Each output channel can potentially represent a separate task. We train models with up to four output channels: building semantic segmentation, roads semantic segmentation, building centroids, and super-resolution grayscale image, as shown in Figure 4. We use the building centroid predictions with one extra step to calculate building counts. We predict a super-resolution grayscale image as one of the output channels as this helps with image registration during training and evaluation.
@@ 4 Dataset
The training split of the dataset consists of about 10 million stacks of Sentinel-2 imagery sampled globally. Examples are sampled randomly at uniform first and then subsampled by building presence using the high-resolution teacher model so that approximately 90% of the examples have buildings. The geographical distribution of examples is shown in Figure 8. A consequence of our sampling strategy is that the distribution of our dataset is roughly correlated with population density. The validation split consists of 1165 Sentinel-2 imagery stacks, which are all within the continent of Africa and chosen to include a range of urban and rural examples of different densities, as well as settings of humanitarian significance such as refugee settlements. Training examples close to validation examples (within some radius) are discarded. For simplicity we do not perform any sampling based on roads. Often roads tend to be present in the vicinity of buildings, but this is not always true.
@@ 4.1 Teacher labels
Figure 7: Geographical distributions of high-resolution, human labelled data on which the teacher model was trained, and human-labelled evaluation data on which we report all evaluation metrics. Counts per level 5 S2 cell.
Figure 8: Geographical distribution of teacher-labelled training data used to train the student model. Counts per level 5 S2 cell.
@@ 4.2 Registration
Neither low-resolution Sentinel-2 images nor the high-resolution image are registered to any reference frame, and as such, both the input imagery stack as well as the labels are potentially misaligned relative to each other. Given the resolution of Sentinel-2 imagery, misalignments of a few pixels in image space amount to tens of meters on the ground. We rely on the model being able to implicitly align the low-resolution input frames and to facilitate such implicit alignment, inspired by [10], we tried pairing each input frame with the frame closest in time to the high-resolution image used to derive labels, i.e. the 17th frame. The model output however can still be misaligned relative to the labels, and in order to not penalise the model for that we manually align labels to model output during loss computation, as described in Section 6. We also do similar alignment in evaluation.
@@ 4.4 Orthorectification
Label transfer from high resolution to low resolution works well only if both the satellites have the same viewing angle or both the imagery collections are orthorectified. The high-resolution imagery collection we use is however not orthorectified (see example in Figure 10) and as such, labels generated using such imagery do not always align with Sentinel-2 imagery. Particularly problematic are areas with tall buildings where building roofs (which is what our teacher model is trained to detect) appear in different positions depending on the satellite viewing angle. Orthorectification of satellite imagery however requires an accurate elevation model – something that is not available globally. In the absence of orthorectification, labels tends to be more accurate for rural areas than for urban areas with taller buildings.
@@ 4.5 Sentinel-2 data processing
We feed in metadata associated with each frame to the model using a simple approach where scalar value for each metadata item is broadcast to image spatial dimensions and then appended to the image in the channel dimension. As such, each scalar metadata value adds an additional channel to the input. The following metadata are used for each frame: normalized time relative to the 17th time frame, mean incidence azimuth angle, mean incidence zenith angle, mean solar azimuth angle, mean solar zenith angle, latitude, and longitude. The time of image acquisition is normalised by dividing the time duration relative to the 17th frame by ten years in seconds, and scale the other features to the [0, 1] range. We do not perform any data augmentation other than random cropping, which we found is sufficient to prevent overfitting due to the very large dataset size.
@@ 5.1 Encoder
We employ the HRNet [15] architecture (see Figure 11) as an encoder without making any significant modifications, except for the root block. The original root block in HRNet downsamples the input image by a factor of 4, which we find to be too aggressive for LR input images that are already quite low resolution. To address this, we remove the 4x downsampling by reducing stride from 2 to 1 on the two 3x3 convolutions in the first block of HRNet. As a result, the encoder outputs features at the same spatial resolution as the input. We pre-train the HRNet encoder on ImageNet. To adapt filter weights of the first convolution trained on 3 channel RGB ImageNet input to more channels, for each filter we take the mean of the weights across channels and replicate the mean across the target channel dimensions (=13+7, corresponding to all bands in a Sentinel-2 image and all metadata we pass in).
@@ 5.3 Decoder
The decoder module in our model consists of a number of decoder blocks that each upsample the input by a factor of 2. We use the same decoder block as in our previous technical report [14] which consists of x2 (batch normalization, ReLU, convolution) followed by another (batch normalization, ReLU), fusion with residual connection to the input and finally an upconvolution, as illustrated in Figure 14. We use transposed convolution or deconvolution for up convolution. The fused encoding is passed through 3 decoder blocks with widths 360, 180 and 90 respectively. Finally the upsampled features are passed through a (convolution, batch normalization, ReLU) block before being passed through a final convolution that outputs the desired number of output channels. There is one output channel for each task: building segmentation, road segmentation, centroid detection and image super-resolution. Since the input Sentinel-2 frames are the same for each of these tasks, a single model can be trained in a multi task fashion to perform these tasks simultaneously. Apart from reducing cost and time, the different tasks could possibly serve as regularization for the other tasks to produce better results.
@@ 6 Loss function
where $y_{i}$ and $\hat{y_{i}}$ are the label and prediction for pixel $i$ respectively and $\gamma$ is the focal term. Both $y_{i}$ and $\hat{y_{i}}$, originally $\in[0,1]$, are clipped to $[\epsilon,1-\epsilon]$, where $\epsilon=1e-7$, to avoid division by zero errors. The focal term $\gamma$ varies the importance given to hard (misclassified) examples. Following a hyper-parameter sweep we set $\gamma$ to 0.25. We also add $\epsilon$ to KLD before exponentiation by $\gamma$ to guard against an undefined gradient.
@@ 7 Experiments
The numbers between different tables are not necessarily comparable due to slight differences in configurations. Our evaluations focus on the building detection task and we do not report metrics for the road detection and image super-resolution tasks.
@@ 7.3 Pairing schemes
Inspired by HighResNet [10], where each low-resolution frame is paired with the per-pixel median of the stack, we experimented with various pairing schemes. We find that pairing each timeframe with the 17th timeframe (that is closest in time to the teacher label) provides a significant boost in performance over model without pairing. In comparison, model trained with no pairing but with cross-time communication (Section 5.2) performs slightly worse (see Table 1). Surprisingly enough, model trained with both pairing (17th frame) and cross-time communication performs worse than that with just pairing. We also experimented with pairing with either the median (as is done in HighResNet), mean or second darkest timeframe and find none of these pairing schemes to be better than pairing with the 17th frame.
@@ 7.4 Resolution sensitivity analysis
50 85.3
Best Sentinel-2 model 78.3
@@ 8.1 Social and ethical considerations
Method mIoU (Buildings)
Only future 75.8
Only past 76.6
Equal past & future 76.3

### v3

@@ Abstract
Mapping buildings and roads automatically with remote sensing typically requires high-resolution imagery, which is expensive to obtain and often sparsely available. In this work we demonstrate how multiple 10 m resolution Sentinel-2 images can be used to generate 50 cm resolution building and road segmentation masks. This is done by training a ‘student’ model with access to Sentinel-2 images to reproduce the predictions of a ‘teacher’ model which has access to corresponding high-resolution imagery. While the predictions do not have all the fine detail of the teacher model, we find that we are able to retain much of the performance: for building segmentation we achieve 79.0% mIoU, compared to the high-resolution teacher model accuracy of 85.5% mIoU. We also describe two related methods that work on Sentinel-2 imagery: one for counting individual buildings which achieves $R^{2}=0.91$ against true counts and one for predicting building height with 1.5 meter mean absolute error. This work opens up new possibilities for using freely available Sentinel-2 imagery for a range of tasks that previously could only be done with high-resolution satellite imagery.
@@ 1 Introduction
Figure 3: Estimation of the number of buildings in a tile, based on predicting building centroids (left: high resolution image for comparison, centre: Sentinel-2 RGB; right: predicted centroid mask). This method can obtain $R^{2}=0.91$ with respect to true counts even though individual buildings cannot be discerned in the source imagery.
Figure 4: Above-ground object height prediction (left: high resolution image for comparison, centre: Sentinel-2 RGB; right: predicted height mask). This method can predict building height with 1.5 meter mean absolute error.
In our work, we attempt to extend the limit of fine detail that can be recreated from a set of Sentinel-2 images, assessing the quality of building and road presence predictions made at 50 cm resolution. To do this, we use a teacher model with access to 50 cm resolution imagery to create training labels for a large dataset of worldwide imagery. We then train a student model to reconstruct these labels given only a stack of Sentinel-2 images from the corresponding places and times (see Figure 1). We find that we are able to retain much – though not all – of the accuracy of the high-resolution teacher model: our Sentinel-2-based building segmentation has 79.0% mIoU, compared to the high-resolution-based teacher performance of 85.5% mIoU. We found that this accuracy level was equivalent to what could be achieved by a high-resolution model using a single frame of 4 m resolution imagery (see Table 5), though also noting through visual inspection that the segmentation quality sometimes far exceeds that (see Figure 2 and Section 10).
We additionally describe experiments using the same model framework for two additional tasks: counting buildings and estimating building height. Building counts are estimated based on predicting the locations of building centroids. We can obtain $R^{2}=0.91$ compared to true building counts, again capturing much of the performance of the high-resolution teacher model ($R^{2}=0.95$). The predicted counts are surprisingly accurate even when buildings are very small relative to the raw Sentinel-2 resolution, or close together. Some examples are shown in Figures 3 and 16. Height prediction is done using per pixel above-ground object height labels during training. We find that the model is able to learn to predict building height with 1.5 meter mean absolute error. Examples are shown in Figures 4 and 24.
This work extends the range of analysis tasks that can be carried out with freely available Sentinel-2 data. Buildings data produced by this model across Africa, South and Southeast Asia, and Latin America is now publicly available [^1] https://sites.research.google/gr/open-buildings/temporal/.
@@ 3 Training setup
Each output channel can potentially represent a separate task. We train models with up to five output channels: building semantic segmentation, roads semantic segmentation, building centroids, building heights and super-resolution grayscale image, as shown in Figure 5. We use the building centroid predictions with one extra step to calculate building counts. We predict a super-resolution grayscale image as one of the output channels as this helps with image registration during training and evaluation.
@@ 4 Dataset
The training split consists of two datasets: first one with about 10 million examples sampled globally where unorthorectified high-resolution satellite imagery was available and second one with 2.5 million examples sampled where high-resolution orthorectified imagery with detailed DSM (Digital Surface Model) and DTM (Digital Terrain Model) were available i.e. mostly in North America, Europe and Japan. For the first dataset, the examples are sampled randomly at uniform first and then subsampled by building presence using the high-resolution teacher model so that approximately 90% of the examples have buildings, therefore as a consequence the distribution is roughly correlated with population density. The geographical distribution of examples is shown in Figure 8.
There are four validation datasets with human-derived building instance labels: ”Africa” with 1158 examples, ”Africa hard” with 945 examples, ”South-Southeast Asia” with 1620 examples, and ”Latin America” with 1631 examples. Furthermore, a separate validation dataset with above-ground object height labels (see Section 4.2) but no human-derived building instance labels that consists of 1029 examples. The Africa dataset has examples which are all within the continent and chosen to include a range of urban and rural examples of different densities, as well as settings of humanitarian significance such as refugee settlements. ”Africa hard” contains examples sampled from Africa where earlier high-resolution models had poor performance. The South-Southeast Asia and Latin America validation splits consist of examples where 75% were sampled randomly across the regions and the other 25% were sampled randomly while making sure environments with diverse building counts, building sizes and total building area were well represented. Training examples close to validation examples (within some radius) are discarded. For simplicity we do not perform any sampling based on roads. Often roads tend to be present in the vicinity of buildings, but this is not always true.
@@ 4.1 Teacher labels
(a) Teacher training: human labels (400K)
(b) Student training: teacher labels (10M)
(c) Student training: heights and teacher labels (2.5M)
(d) Validation: heights and teacher labels (1K)
(e) Val.: Africa (1.2K)
(f) Val.: Africa hard (1K)
(g) Val.: South-Southeast Asia (1.6K)
(h) Val.: Latin America (1.6K)
Figure 8: Geographical distribution of all datasets used in this work. The teacher model was trained on the human-labelled high-resolution unorthorectified dataset in (8(a)). The student model was trained on teacher-labelled unorthorectified dataset (8(b)) and teacher-labelled orthorectified dataset with height labels (8(c)). Furthermore, the student model was evaluated on datasets (8(d))-(8(h)). The datasets (8(e))-(8(h)) are unorthorectified, human-labelled, and without height labels. Counts per level 5 S2 cell. The total number of examples is provided in parentheses.
4.2 Above-ground object height labels
Figure 9: Above-ground object height label example.
Approximately 20% of the examples in the datasets were sampled in locations where we had access to a detailed DSM (Digital Surface Model) and DTM (Digital Terrain Model) matching time-wise the orthorectified high-resolution image. By subtracting the DTM label from the DSM label we obtained above-ground object height labels that for a given pixel denote the height in meters of objects such as buildings, bridges and trees. Before the height labels were generated the DTM label was smoothed out to a consistent value for all pixels within a single building instance. As a consequence of how labels are generated the model is, among others, able to predict vegetation height. However, in this work we only focus on the building height prediction aspect, therefore for height prediction evaluation we ignore the non-building pixels using the teacher labels (human-derived building labels were unavailable). An example of a above-ground object height label is shown in Figure 9.
@@ 4.3 Registration
Neither low-resolution Sentinel-2 images nor the high-resolution image are registered to any reference frame, and as such, both the input imagery stack as well as the labels are potentially misaligned relative to each other. Given the resolution of Sentinel-2 imagery, misalignments of a few pixels in image space amount to tens of meters on the ground. We rely on the model being able to implicitly align the low-resolution input frames. However, model output can still be misaligned relative to the labels, and in order to not penalise the model for that we manually align labels to model output during loss computation, as described in Section 6. We also do similar alignment in evaluation.
@@ 4.5 Orthorectification
Label transfer from high resolution to low resolution works well only if both the satellites have the same viewing angle or both the imagery collections are orthorectified. 80% of the high-resolution imagery we use is however not orthorectified (see example in Figure 11) and as such, labels generated using such imagery do not always align with Sentinel-2 imagery. Particularly problematic are areas with tall buildings where building roofs (which is what our teacher model is trained to detect) appear in different positions depending on the satellite viewing angle. Orthorectification of satellite imagery however requires an accurate elevation model – something that is not available globally (see Figure 8(c)). In the absence of orthorectification, we pass the high-resolution satellite image incidence angle metadata on input to the model. See Section 7.8 for more details.
@@ 4.6 Sentinel-2 data processing
We feed in metadata associated with each frame to the model using a simple approach where scalar value for each metadata item is broadcast to image spatial dimensions and then appended to the image in the channel dimension. As such, each scalar metadata value adds an additional channel to the input. The following metadata are used for each frame: normalized time relative to the 17th time frame, mean incidence azimuth angle, mean incidence zenith angle, mean solar azimuth angle, mean solar zenith angle, latitude, longitude, mean high-resolution satellite image incidence azimuth angle and zenith angle (the last 4 are the same for each frame). The time of image acquisition is normalised by dividing the time duration relative to the 17th frame by ten years in seconds, and scale the other features to the [0, 1] range. We do not perform any data augmentation other than random cropping, which we found is sufficient to prevent overfitting due to the very large dataset size. Mean high-resolution satellite image incidence azimuth angle and zenith angle are set to $0^{\circ}$ during inference.
@@ 5.1 Encoder
We employ the HRNet [15] architecture (see Figure 12) as an encoder without making any significant modifications, except for the root block. The original root block in HRNet downsamples the input image by a factor of 4, which we find to be too aggressive for LR input images that are already quite low resolution. To address this, we remove the 4x downsampling by reducing stride from 2 to 1 on the two 3x3 convolutions in the first block of HRNet. As a result, the encoder outputs features at the same spatial resolution as the input. We pre-train the HRNet encoder on ImageNet. To adapt filter weights of the first convolution trained on 3 channel RGB ImageNet input to more channels, for each filter we take the mean of the weights across channels and replicate the mean across the target channel dimensions (=13+9, corresponding to all bands in a Sentinel-2 image and all metadata we pass in).
@@ 5.3 Decoder
The decoder module in our model consists of a number of decoder blocks that each upsample the input by a factor of 2. We use the same decoder block as in our previous technical report [14] which consists of x2 (batch normalization, ReLU, convolution) followed by another (batch normalization, ReLU), fusion with residual connection to the input and finally an upconvolution, as illustrated in Figure 15. We use transposed convolution or deconvolution for up convolution. The fused encoding is passed through 3 decoder blocks with widths 360, 180 and 90 respectively. Finally the upsampled features are passed through a (convolution, batch normalization, ReLU) block before being passed through a final convolution that outputs the desired number of output channels. There is one output channel for each task: building segmentation, road segmentation, building centroid detection, building height prediction and image super-resolution. Since the input Sentinel-2 frames are the same for each of these tasks, a single model can be trained in a multi task fashion to perform these tasks simultaneously. Apart from reducing cost and time, the different tasks could possibly serve as regularization for the other tasks to produce better results.
@@ 6 Loss function
where $y_{i}$ and $\hat{y_{i}}$ are the label and prediction for pixel $i$ respectively and $\gamma$ is the focal term. Both $y_{i}$ and $\hat{y_{i}}$, originally $\in[0,1]$, are clipped to $[\epsilon,1-\epsilon]$, where $\epsilon=1e-7$, to avoid division by zero errors. The focal term $\gamma$ varies the importance given to hard (misclassified) examples. Following a hyper-parameter sweep we set $\gamma$ to 0.25. We also add $\epsilon$ to KLD before exponentiation by $\gamma$ to guard against an undefined gradient. For examples where the above-ground object height label is missing we zero out the loss term for the height task.
@@ 7 Experiments
The numbers between different tables are not necessarily comparable due to slight differences in configurations. Our evaluations focus on the building detection task and we do not report metrics for the road detection and image super-resolution tasks. Unless otherwise noted evaluation numbers are reported for the Africa validation dataset.
@@ 7.3 Pairing schemes
Inspired by HighResNet [10], where each low-resolution frame is paired with the per-pixel median of the stack, we experimented with various pairing schemes. We find that pairing each timeframe with the 17th timeframe (that is closest in time to the teacher label) provides a significant boost in performance over model without pairing. In comparison, model trained with no pairing but with cross-time communication (Section 5.2) performs slightly worse (see Table 1). Surprisingly enough, model trained with both pairing (17th frame) and cross-time communication performs worse than that with just pairing. We also experimented with pairing with either the median (as is done in HighResNet), mean or second darkest timeframe and find none of these pairing schemes to be better than pairing with the 17th frame. Ultimately, we used cross-time information fusion and no pairing scheme, based on the observation that this method appeared more robust in cases where the 17th frame was cloudy.
@@ 7.4 Resolution sensitivity analysis
50 85.5
Best Sentinel-2 model 79.0
@@ 7.6 Number and position of Sentinel-2 timeframes
Method mIoU (Buildings)
Only future 75.8
Only past 76.6
Equal past & future 76.3
@@ 7.7 Building counts via centroid prediction
7.8 Incidence angle metadata for high-resolution imagery
As mentioned in Section 4.5, 80% of the high-resolution images we use are not orthorectified. This is a problem especially for tall buildings as the student model is not able to guess the position of the roofs in the high-resolution image from which the teacher label was generated and hence tends to produce blurry output in those cases. To mitigate this issue during training we pass the high-resolution satellite image incidence angle metadata on input to the student model, so that it can learn to modify its output to match the teacher label. During inference we set the incidence angle to $0^{\circ}$ so that the model predicts orthorectified output. See Figure 19 and Figure 20 for an example of visual quality improvement in model output.
Figure 19: Output from a model trained without high-resolution incidence metadata in Manhattan.
Figure 20: Output from a model trained with high-resolution incidence metadata in Manhattan.
7.9 Building height prediction
As described in Section 4.2, 20% of examples in the datasets have per pixel above-ground object height labels that we used to train the model to solve the auxiliary task of building height prediction (capped at 100 meters). The best model was able to achieve mean absolute error of 1.5 meters on the building instance height validation dataset. Additional evaluation results for height prediction can be seen in Table 9. Further research is necessary to understand what cues the model uses for height prediction; for example this may be related to shadow length, or a parallax effect due to Sentinel-2 channels being captured at slightly different times as the satellite passes overhead.
Table 9: Mean and percentiles of building instance height absolute error (AE) of the best model on the height prediction validation dataset.
Building height range Mean AE 50th pct. AE 90th pct. AE 95th pct. AE 99th pct. AE
(0 m; 5 m) 1.01 m 0.76 m 2.15 m 2.71 m 4.54 m
[5 m; 20 m) 1.63 m 1.14 m 3.71 m 4.86 m 7.87 m
[20 m; 100 m] 3.64 m 2.53 m 7.95 m 11.00 m 17.95 m
(0 m; 100 m] 1.50 m 0.99 m 3.35 m 4.66 m 8.48 m
7.10 Performance across different validation datasets
We evaluated the best model across all validation datasets on building segmentation using mean Intersection over Union (mIoU) metric, on building counting using the coefficient of determination ($R^{2}$) metric and on total built-up area prediction using relative error metric. For comparison, we included mIoU of the teacher model on the high-resolution images. Results are summarized in Table 10 and the score thresholds used for binarization that maximize the metrics are in Table 11. The model performs well across all validation datasets with notable exception of building counting on the ”Africa hard” dataset. The gap in performance between the Sentinel-2 model and the high-resolution model varies 3.3-6.5 mIoU and is the largest on the Africa validation dataset. Moreover, the score thresholds are similar across all validation datasets.
Table 10: The best model evaluated on all building segmentation validation datasets.
Dataset Teacher model mIoU mIoU Building count $R^{2}$ Built-up area error
Africa 85.5 79.0 0.912 0.79%
Africa hard 85.2 80.0 0.837 0.15%
South-Southeast Asia 87.2 82.7 0.936 0.79%
Latin America 87.7 84.4 0.887 0.18%
Table 11: Score thresholds used for binarization that maximize the metrics for the best model.
Dataset Best mIoU score threshold Best built-up area error score threshold
Africa 0.34 0.40
Africa hard 0.37 0.41
South-Southeast Asia 0.34 0.41
Latin America 0.37 0.45